# Fast sampling and model selection for Bayesian mixture models

### v1

@@ IV.1 Latent class analysis
Latent class analysis (LCA) uses mixture modeling to classify patterns of responses to test and survey data, or other data with categorical variables Goodman74 ; McCutcheon87 ; BR93 ; LLSL18 . In this paper we discuss LCA in the language of survey data, but the methods we describe are broadly applicable to any categorical data.
@@
Consider a survey in which $N$ respondents are divided into $k$ components, also called classes in this context, and give responses to $Q$ questions. Let $\theta_{rqx}$ be the probability that a respondent in class $r$ answers question $q$ with response $x$. Each response is considered independent, conditioned on the class assignments and the response probabilities. If $x_{iq}$ is the answer given by respondent $i$ to question $q$, then we can write the complete data likelihood as
@@
where $m_{rqx}$ is the number of respondents in class $r$ who answered question $q$ with response $x$. One normally assumes a symmetric Dirichlet prior on the $\theta_{rqx}$ for each class/question combination, with some concentration parameter $\alpha$, which gives
$P(x,\theta|k,z,\alpha)=\prod_{rq}{\prod_{x=1}^{k_{q}}\theta_{rqx}^{m_{rqx}+\alpha-1}\over\mathop{\textrm{B}}(\alpha,\alpha,\ldots)},$ (17)
where $k_{q}$ is the number of distinct possible answers to question $q$ and $\mathop{\textrm{B}}(\alpha,\alpha,\ldots)$ is the multivariate beta function. Most commonly one makes the uniform choice $\alpha=1$, which is often a good assumption, although there may also be reason to believe that the distribution of answers across classes will be imbalanced, in which cases a smaller value of $\alpha$ may be appropriate. The choice $\alpha=\frac{1}{2}$, for instance, which is the Jeffreys prior for the Dirichlet distribution, is not uncommon. For the calculations performed in this paper we use $\alpha=1$, but we present the derivations for general $\alpha$ and our code implements the general $\alpha$ case.
Integrating over the $\theta$ parameters, we now have
$P(x|k,z,\alpha)$ $=\int P(x,\theta|k,z,\alpha)\>\mathrm{d}\theta$
$=\prod_{rq}{1\over\mathop{\textrm{B}}(\alpha,\alpha,\ldots)}\int\prod_{x=1}^{k_{q}}\theta_{rqx}^{m_{rqx}+\alpha-1}\>\mathrm{d}\theta$
$=\prod_{rq}{\Gamma(\alpha k_{q})\over\Gamma(n_{r}+\alpha k_{q})}\prod_{x=1}^{k_{q}}{\Gamma(m_{rqx}+\alpha)\over\Gamma(\alpha)},$ (18)
where $\Gamma(x)$ is the Euler gamma function and we have made use of the fact that $\sum_{x}m_{rqx}=n_{r}$ for all $q$. Combining this likelihood with Eq. (5) now allows us to write a complete posterior $P(k,z|x,\alpha)=P(x|k,z,\alpha)P(z|k)P(k)/P(x)$, up to the unknown constant $P(x)$.
For the prior $P(k)$ on the number of classes we assume the uniform choice $P(k)=1/N$, which then gives simple expressions for the weights in our Monte Carlo algorithm. For the conventional collapsed Gibbs sampler, our algorithm 1, Eqs. (10) and (11) become
$w_{\mu}$ $={N-k\over k}(n_{s}+1)\prod_{q}{m_{sqx_{iq}}+\alpha\over n_{s}+\alpha k_{q}},$ (19)
$w_{\mu}$ $=(k+1)\prod_{q}{1\over k_{q}},$ (20)
respectively, and for the enhanced sampler of algorithm 2, Eqs. (13) and (14) become
$w_{\mu}$ $={N-k\over k}\prod_{q}{m_{sqx_{iq}}+\alpha\over n_{s}+\alpha k_{q}},$ (21)
$w_{\mu}$ $=k\prod_{q}{1\over k_{q}}.$ (22)
Algorithm 2 is the primary algorithm that we focus on in this paper, since it has generally superior performance. We measure performance in terms of “sweeps,” where one sweep means $N$ successive Monte Carlo steps. The length of the run needed to adequately sample class assignments varies from one data set to another, but for most of the data sets we study we find good results with 2500 Monte Carlo sweeps for burn-in followed by 25 000 sweeps for actual sampling, with one sample taken at the end of each sweep. More generally, one should run for long enough to ensure adequate sampling of the space of class assignments. A simple way to do this is to perform multiple runs and confirm that the results are consistent across runs to the desired accuracy. By these standards the 25 000-sweep runs used in our calculations appear more than adequate in most cases, although for one data set we found it necessary to run for longer—see Section IV.4.
A discussion of some additional technical aspects of our two algorithms is given in Appendix A, and a complete implementation is available at https://umich.edu/~mejn/lca for download.
Running time ns/step Acceptance ratio Correlation time
Data set $N$ $Q$ Alg. 1 Alg. 2 Alg. 1 Alg. 2 Alg. 1 Alg. 2 Alg. 1 Alg. 2
Synthetic data 1000 8 8.7 9.2 85 90 23.3% 33.4% 155.0 27.8
Opinion poll 1566 32 30.1 32.8 190 210 14.7% 14.5% 7.8 7.5
Alzheimer’s symptoms 240 6 0.7 0.8 30 30 27.9% 32.9% 162.6 53.6
Teenage behavior 6504 6 32.0 33.4 50 50 29.7% 37.7% 2837.5 1498.3
50 states 50 12 0.2 0.2 40 45 8.7% 14.6% 22.0 9.0
Online dating 59 946 14 2150 2140 360 355 50.2% 47.4% 73.7 132.6
Table 1: Performance measures for the two algorithms introduced in this paper, applied to latent class analysis and various data sets. Columns are: number $N$ of observations (respondents), number of questions $Q$, running time in seconds for each algorithm, nanoseconds per Monte Carlo step, acceptance ratio, and integrated correlation time measured in sweeps. Running times represent CPU seconds for 100 000 sweeps (after burn-in), measured on the author’s laptop.
IV.2 Performance measures
Table 1 gives performance figures for applications of our two algorithms to a selection of data sets typical of latent class analyses. The individual data sets are discussed in more detail in the following sections and in Appendix B.
For each data set the table lists total running time and running time per Monte Carlo step. The code is fast, performing millions of Monte Carlo steps per second, and total running times for most calculations are a few seconds. Running time and time per step are typically slightly shorter for algorithm 1 than for algorithm 2, but this is deceptive. Some Monte Carlo steps leave the state unchanged, as mentioned previously. Such null moves are faster, because less work is involved, but they are also essentially wasted computer time, since they don’t do anything. It turns out that algorithm 1 performs more null moves than algorithm 2, which makes it faster per step, but at the expense of actually achieving less. The magnitude of this effect can be quantified by the acceptance ratio, the fraction of Monte Carlo steps that actually change the state of the system. Acceptance ratios for our two algorithms are also given in Table 1. Values range from about 8% to over 50%, which are considered good figures in this context: acceptance ratios for naive Metropolis-Hastings Monte Carlo methods can be as low as 1% or less. Moreover, the acceptance ratio is often higher for algorithm 2 than for algorithm 1, which results in improved performance—fewer Monte Carlo steps need be performed to generate a given number of accepted moves.
Even these numbers do not tell the full story, however, because accepted Monte Carlo moves do not necessarily translate into independent samples. The number of moves needed to generate an independent sample can be quantified by calculating the autocorrelation time of the samples drawn. In this paper we compute the autocorrelation from the power spectrum of the posterior probability and then calculate the integrated correlation time NB99 . Figures for the resulting correlation times, measured in sweeps, are given in Table 1 for both our algorithms. This is where we see a significant difference between the algorithms, the correlation time being substantially shorter for our algorithm 2 in most cases, by up to a factor of five or more, which improves performance greatly—many fewer sweeps are necessary to generate a given number of independent samples.
Figure 1: (a) Density plot of the consensus matrix for a single run of the Monte Carlo algorithm on a synthetic data set with 1000 respondents, 10 questions, and 8 classes. The classes are clearly visible as the boxes down the diagonal of the matrix. (b) The inferred posterior distribution over the number $k$ of classes, which peaks at the correct value of 8, but gives significant weight to $k=7$ and $k=9$ as well.
Two of our data sets illustrate interesting special cases that are worth mentioning. The first is the data set on problem behaviors among teenagers (see Section IV.4), which has a much longer correlation time than any of the others, by a factor of 20 or so. It is not clear what makes mixing so much slower for this data set, but the difference has practical repercussions—as reported in Section IV.4, we find it necessary to run for considerably longer to generate consistent results. For all of the other data sets, the correlation time is around 50 sweeps at most, which means that a run of 25 000 sweeps, as used in most of our calculations, draws about 500 independent samples, enough to estimate expected values, such as number of classes and coincidence rates, to within a few percent. The speed of the algorithm, however, means that longer runs are also within easy reach. For the teenage behaviors data set we ran for 250 000 sweeps, which takes about 90 seconds.
The other notable data set is the online dating one, for which the acceptance ratios of the two algorithms are nearly identical and the traditional Gibbs sampler actually mixes faster. Overall we recommend using algorithm 2 for its generally better efficiency, but it is worth bearing in mind that occasionally the traditional sampler of algorithm 1 can be (slightly) faster.
It would be possible to speed up our calculations further by parallelization. The bulk of the computational effort goes into calculating the weights $w_{\mu}$ of Eqs. (13) and (14), which has to be done anew for every Monte Carlo step. The individual weights however are not dependent and can be calculated in any order, which makes the algorithm trivially parallelizable and an ideal candidate for a multithreaded implementation.
IV.3 Synthetic tests
@@
In the next few sections we give details on applications of our methods to our various data sets, focusing mainly on algorithm 2 and demonstrating the algorithm’s ability to directly infer the number of classes in the data as well as the class assignments. For our first example, we test the method against synthetic (i.e., computer-generated) data, generated from the same model of Section II used in the Monte Carlo inference procedure. These calculations can be thought of as consistency tests and we focus particularly on whether we are able to correctly recover the number of classes present in the data.
The synthetic data sets we generate have $N=1000$ respondents, $Q=10$ questions, $k_{q}=4$ possible answers to each question, a Dirichlet concentration parameter of $\alpha=1$, and varying numbers of classes. Figure 1 shows results from a typical run with eight planted classes. The running time was about 3 seconds. Panel (a) shows a visualization of the consensus matrix (also sometimes called the association matrix), the matrix whose elements are the coincidence rates (the fraction of samples in which a pair of respondents are in the same class together). If we draw $T$ samples of the class assignments, then the estimated consensus matrix is the $N\times N$ matrix with elements
$C_{ij}={1\over T}\sum_{t=1}^{T}\mathbbm{1}_{z^{(t)}_{i}=z^{(t)}_{j}},$ (23)
where $z^{(t)}$ are the class assignments in sample $t$ and $\mathbbm{1}_{x}$ is the indicator function which is 1 if $x$ is true and 0 otherwise. It is clear from the figure that the algorithm has successfully identified the eight classes in the data, as indicated by the prominent blocks along the diagonal of the consensus matrix, which correspond to sets of respondents that are grouped together most of the time. (Without loss of generality, we create the classes as contiguous blocks of respondents, so that their structure is visually obvious in plots such as Fig. 1a.) Two classes, the right-most two blocks in the figure, are weaker, and more weakly distinguished from one another, presumably because the values of the randomly chosen response probabilities $\theta_{rqx}$ happen to be quite similar, so there is relatively little signal of class membership in the simulated data. This is a realistic situation—it could certainly occur in real-world data sets and we would hope that the algorithm would signal some uncertainty about the number of classes $k$ in such cases.
Figure 1b shows the posterior distribution over the number of classes, which is simply a histogram of the sampled values of $k$. We see that the algorithm has again been successful—the most likely (MAP) value is the correct $k=8$, although $k=7$ is also competitive, as we might expect, as is $k=9$ to a lesser extent. This is typical of results from these methods and should be considered a strong point of the approach. There is genuine uncertainty inherent in the data about the number of classes, which the Bayesian approach reveals in a way that other methods cannot.
@@
For a more systematic test of the method, we repeat these calculations on a large number of synthetic data sets of the same size and parameters, but with varying numbers of classes $k$ from 2 up to 20. For each value of $k$ we generate 1000 data sets and use our Monte Carlo algorithm to compute the MAP estimate of $k$ for each one, which is simply the most commonly occurring value of $k$ among all samples drawn. Figure 2 shows the results plotted against the ground-truth values used to generate the data. In this figure the violin plots represent the variation in $k$ across all 1000 repetitions and the circles represent the median values. If the algorithm has learned the number of classes correctly the data should fall on the diagonal dashed line and, as the figure shows, the median results are perfect in this case. There is, however, considerable variation about the median, which occurs because our synthetic data sets can be genuinely ambiguous, as we saw in Fig. 1. Again this is one of the strengths of the approach, that it clearly reveals and quantifies uncertainties where present.
@@
IV.4 Real-world data
In this section we apply our methods to a number of real-world data sets, drawn from different spheres. The data sets are described in detail in Appendix B.
IV.4.1 CBS/New York Times opinion poll
@@
We perform a single Monte Carlo run on this data set of 2500 sweeps for burn-in followed by 25 000 sweeps to take samples, which produces the results shown in Fig. 3. As we can see, the method is confident in this case that there are eight classes in the data. The classes can be seen clearly in the plot of the consensus matrix, although one class, represented by the second-to-last block in the matrix, is very small, with only five members, making it more difficult to see.
IV.4.2 Diagnostic survey for Alzheimer’s disease
Our second example is a clinical one: we analyze data from Walsh Moran04 ; Walsh06 describing observations of potential cognitive decline in 240 patients at a memory clinic in Ireland. The survey instrument asked primary caregivers about the presence or absence in the patients of six symptoms commonly associated with Alzheimer’s disease: activity disturbance, affective disorder, aggression, agitation, diurnal rhythm disturbance, and hallucinations. Figure 4 shows the consensus matrix and distribution of the number of classes for these data and, as we can see, there are most likely two groups in the data, although the division between the classes is less clear than for the previous examples and the posterior distribution on $k$ allows for values up to $k=7$. Walsh also found two classes in his work, corresponding to patients with and without substantial symptoms, although he comments that there is potentially interesting additional structure in the three-class division, which carries a nontrivial posterior probability in our analysis (Fig. 4b). White et al. WWM16 , analyzing the same data using a Monte Carlo method based on the allocation sampler of NF07 , reached similar conclusions, although their calculation took over 3 minutes, where ours takes a fraction of a second.
@@ IV.4.2 Diagnostic survey for Alzheimer’s disease
The roughly block-diagonal form of the consensus matrix seen here and in our other examples implies that the matrix is approximately of low rank, which suggests a possible further analysis: we could look at the leading eigenvectors of the matrix in a manner akin to principal components analysis or correspondence analysis. We do this in Fig. 5 and the figure does a good job of revealing structure and clustering in the data, which in turn suggests a natural method of consensus clustering in which we group the points on such a plot using, for example, $k$-means clustering, with the number of classes chosen to match the MAP estimate from the Monte Carlo results (which is two in this case). This approach has been studied recently by Zhou et al. ZDA23 , who show that it has provably good performance under a simple model of label perturbation.
@@
IV.4.3 Problem behaviors among teenagers
Our third example application is to a large behavioral data set and illustrates a case in which our method gives different answers from previous approaches. We examine a data set presented by Li et al. LLSL18 , which describes survey results on the presence or absence of six problem behaviors in 6504 US teenagers. A previous analysis of a subset of these data by Collins and Lanza CL10 found four latent classes. Figure 6 shows the posterior distribution of the number of classes returned by our Monte Carlo method and, as we can see, four classes is a possible fit in this case, but the MAP estimate is five, and six is also more likely than four. We find a longer run is necessary to achieve consistent results with this data set. We ran for 250 000 Monte Carlo sweeps, plus 25 000 for burn-in. Nonetheless the calculation is fast: Li et al. report that their Gibbs sampling calculation took about half an hour; ours takes less than 90 seconds on roughly comparable hardware.
IV.4.4 Geographic and socio-political traits of the 50 states
Our fourth example is an application to non-survey data. This data set is a geographic one that focuses on the 50 United States and records 12 characteristics for each one, including their position on a range of hot-button issues such as abortion access, gun laws, and the death penalty, as well as more neutral variables such as weather, educational attainment, and whether the state has a professional American football team. The full list of variables is given in Table 2.
On this comparatively small data set our method runs very quickly, completing 25 000 Monte Carlo sweeps in under a second and finding three clear classes, as shown in Fig. 7a. The classes correspond to recognized political divisions among the states: there is a group of left-leaning states (so-called “blue states”), a group of right-leaning ones (“red states”), and a group whose positions on the issues are a mix of left and right. Figure 7b shows the division on a map.
@@ IV.4.4 Geographic and socio-political traits of the 50 states
College education Fraction with a bachelor’s degree
@@
Gun laws Firearm licensing rules
Medicaid Expanded Medicaid under ACA
@@
Same-sex marriage State law permits same-sex marriage
@@
Vote 2024 Winner of presidential vote in 2024
Table 2: Summary of the 12 variables used in the 50-states data set. Further details are given in Appendix B.
@@
This data set provides an opportunity to showcase another application of our method, to variable selection—the identification of variables that are, or are not, particularly informative about class membership RD06 ; DR10 ; GHM11 ; WWM16 . In the present case, for instance, where the classes appear to correspond to political persuasion, one might imagine that the “Temperature” variable, which measures annual average temperature, would not be particularly informative. There are a number of methods for performing variable selection within the latent class analysis paradigm. One approach that tackles the problem head-on is to calculate the mutual information between the responses and the class assignments found by the algorithm RHP22 . Mutual information is precisely a measure of how much one random variable tells us about another, so it directly addresses the question of how much the responses tell us about class membership. For two general variables $x$ and $y$ the mutual information $I$ is defined by
@@
where $P(x)$, $P(y)$, and $P(x,y)$ are the marginal and joint distributions of the variables. In our case, where the variables are the responses $x$ to a question $q$ and assignments to classes $r$, we can calculate these distributions from
@@
IV.4.5 Online dating
For our final example, we demonstrate an application of the method to a much larger data set, which comes from the online dating service OKCupid. Upon creating accounts on this service, users are asked a set of questions about themselves, which the service uses to match them with potential partners. Our data set consists of answers from about 60 000 users to 14 of these questions, which ask about gender, sexual orientation, relationship status, income, offspring, body type, diet, drinking and smoking habits, drug use, zodiac sign, religion, and whether they like cats and dogs.
The analysis of these data takes considerably longer than any of the previous ones, since one sweep now corresponds to 60 000 Monte Carlo steps and hence takes longer to complete. 25 000 sweeps, for example, takes about nine minutes on the author’s laptop. The consensus matrix is too large to draw meaningfully on the page, but Fig. 9 (main panel) shows the posterior distribution of the number of classes and, as we can see, this data set has a much larger number of classes than any of our others, with the distribution strongly peaked around a MAP estimate of $k=31$ classes. Inset in the figure we show the mutual information scores for the 14 questions. As before, these indicate which answers are most informative about class assignment. The most informative answers in this case are (perhaps unsurprisingly) gender and (more surprisingly) whether the respondents like cats and dogs. Apparently the world really is divided between cat people and dog people. High mutual information scores also go to the answers on smoking, drinking, and drugs, perhaps because these three tend to be correlated with one another, and to body type. At the bottom of the list are zodiac sign, which is perhaps not surprising, but also status (single, married, etc.) and sexual orientation. The latter seem more surprising, but we should bear in mind that their low mutual information scores merely indicate that they are not correlated with people’s answers to the other questions. They are saying, for example, that being straight or gay doesn’t affect whether you drink or like dogs.
V Conclusions
In this paper we have described two Monte Carlo algorithms for sampling from the integrated posterior distribution of mixture models. Both algorithms directly sample component assignments and the number of components, providing a way to perform both clustering and model selection in a single calculation, without the need for additional model selection steps. The second of our algorithms, which employs rejection-free sampling from the prior over component assignments, shows particularly good mixing times and acceptance rates.
We have demonstrated our methods with applications to latent class analysis on various data sets, including extensive tests on synthetic data and a selection of real-world examples. We find that the method is consistently able to infer the correct number of classes in the synthetic tests and provides useful results for a range of other tasks, including class assignment, consensus clustering, and variable selection. The method is significantly faster than competing Bayesian methods, with typical runs taking a few seconds to complete.
Acknowledgements.
The author thanks Elizabeth Bruch, Carrie Ferrario, Max Jerdee, and Alec Kirkley for useful conversations. This work was supported in part by the US National Science Foundation under grant DMS–2404617.
Appendix A Proof of correctness for the Monte Carlo algorithm
In this appendix we prove that the distribution of samples drawn by our algorithm 2 (Section III.2, the main algorithm we employ in this paper) converges to the integrated posterior distribution of the corresponding mixture model. We also sketch a proof for algorithm 1, which follows similar lines.

### v2

@@ 4.3 Latent class analysis
For a more substantial test of our algorithm, we turn to latent class analysis (LCA), the mixture modeling of categorical data, such as is used in the analysis of test results or survey data [21, 33, 2, 30]. Here we discuss LCA in the language of survey data, but the methods we describe are broadly applicable to any categorical data.
@@
Consider a survey in which $N$ respondents are divided into $k$ components, also called classes in this context, and give responses to $Q$ multiple-choice questions. Let $\theta_{rqx}$ be the probability that a respondent in class $r$ answers question $q$ with response $x$. If $x_{iq}$ is the answer given by respondent $i$ to question $q$, and assuming independent responses, the complete data likelihood is
@@
where $m_{rqx}$ is the number of respondents in class $r$ who answer question $q$ with response $x$. One normally assumes a symmetric Dirichlet prior on the $\theta_{rqx}$ for each class/question combination, with some concentration parameter $\eta$, which gives
$P(x,\theta|k,z,\eta)=\prod_{rq}{\prod_{x=1}^{k_{q}}\theta_{rqx}^{m_{rqx}+\eta-1}\over\mathrm{B}(\eta,\eta,\ldots)},$ (29)
where $k_{q}$ is the number of distinct possible answers to question $q$ and $\mathrm{B}(\eta,\eta,\ldots)$ is the multivariate beta function again. Integrating over the $\theta$ parameters, we then have
$P(x|k,z,\eta)$ $=\int P(x,\theta|k,z,\eta)\>d\theta=\prod_{rq}{1\over\mathrm{B}(\eta,\eta,\ldots)}\int\prod_{x=1}^{k_{q}}\theta_{rqx}^{m_{rqx}+\eta-1}\>d\theta$
$=\prod_{rq}{\Gamma(\eta k_{q})\over\Gamma(n_{r}+\eta k_{q})}\prod_{x=1}^{k_{q}}{\Gamma(m_{rqx}+\eta)\over\Gamma(\eta)},$ (30)
where we have made use of the fact that $\sum_{x}m_{rqx}=n_{r}$ for all $q$. Combining this expression with Eqs. (14) and (15), the weights for our Monte Carlo algorithm for LCA are
$w_{s}=\prod_{q}{m_{sqx_{iq}}+\eta\over n_{s}+\eta k_{q}},\qquad w_{k+1}={k^{2}\over N-k}{P(k+1)\over P(k)}\prod_{q}{1\over k_{q}},$ (31)
where all quantities, including $k$, denote values before observation $i$ is placed in its new class.
4.3.1 Synthetic tests
@@
As a first test of this algorithm we benchmark it against synthetic data generated from the same model. These calculations can be thought of as consistency tests and we focus particularly on whether we are able to correctly recover the number of classes present in the data.
@@
The synthetic data sets we generate have $N=1000$ respondents, $Q=10$ questions, $k_{q}=4$ possible answers to each question, and a Dirichlet concentration parameter of $\eta=1$. Figure 3 summarizes results from application of our Monte Carlo algorithm to a large number of such data sets with varying numbers of classes $k$ from 2 up to 20. For each value of $k$ we generate 1000 data sets and use our Monte Carlo algorithm to compute the MAP estimate of $k$ for each one, which is simply the most commonly occurring value of $k$ among all samples drawn. Figure 3 shows the results plotted against the ground-truth values used to generate the data. In this figure the violin plots represent the variation in $k$ across all 1000 repetitions and the circles represent the median values. If the algorithm has learned the number of classes correctly the results should fall on the diagonal dashed line and, as the figure shows, the median results are perfect in this case. There is, however, considerable variation about the median. This is typical of results from these methods and should be considered a strong point of the approach. Even for synthetic data there is genuine uncertainty in the data about the number of classes, which the Bayesian approach reveals in a way that other methods cannot.
4.3.2 CBS/New York Times opinion poll
Turning to real-world data sets, we give three examples of applications drawn from different spheres. The data sets are described in detail in Appendix D.
@@ 4.3.2 CBS/New York Times opinion poll
We perform a single Monte Carlo run on this data set of 2500 sweeps for burn-in followed by 25 000 sweeps to take samples, which produces the results shown in Fig. 4. Panel (a) shows the consensus matrix (also sometimes called the association matrix), the matrix whose elements are the coincidence rates (the fraction of samples in which a pair of respondents are in the same class together). In this case the algorithm has identified eight distinct classes in the data, as indicated by the prominent blocks along the diagonal of the consensus matrix, although one class, represented by the second-to-last block in the matrix, is small, with only five members, making it more difficult to see. Figure 4b shows the posterior distribution over the number of classes and we see that the method is confident in this case in its conclusion that there are eight classes in the data (probability 0.74).
4.3.3 Diagnostic survey for Alzheimer’s disease
Our next example is a clinical one: we analyze data from Moran and Walsh (2004; Walsh 60) describing observations of potential cognitive decline in 240 patients at a memory clinic in Ireland. The survey instrument asked primary caregivers about the presence or absence in the patients of six symptoms commonly associated with Alzheimer’s disease: activity disturbance, affective disorder, aggression, agitation, diurnal rhythm disturbance, and hallucinations. Figure 5 shows the consensus matrix and distribution of the number of classes for these data and, as we can see, there are most likely two groups in the data, although the division between the classes is less clear than for the previous example and the posterior distribution on $k$ allows for values up to $k=7$. Walsh [60] also found two classes in his work, corresponding to patients with and without substantial symptoms, although he comments that there is potentially interesting additional structure in the three-class division, which carries a nontrivial posterior probability in our analysis (Fig. 5b). White et al. [62], analyzing the same data using a Monte Carlo method based on the allocation sampler of Nobile and Fearnside [45], reached similar conclusions, although their calculation took over 3 minutes, where ours takes a fraction of a second.
@@ 4.3.3 Diagnostic survey for Alzheimer’s disease
The roughly block-diagonal form of the consensus matrix seen here and in our other examples implies that the matrix is approximately of low rank, which suggests a possible further analysis: we could look at the leading eigenvectors of the matrix in a manner akin to principal component analysis or correspondence analysis. We do this in Fig. 6 and the figure does a good job of revealing structure and clustering in the data, which in turn suggests a natural method of consensus clustering in which we group the points on such a plot using, for example, $k$-means clustering, with the number of classes chosen to match the MAP estimate from the Monte Carlo results (which is two in this case). This approach has been studied by Zhou et al. [63], who show that it has provably good performance under a simple model of label perturbation.
@@
4.3.4 Geographic and socio-political traits of the 50 United States
Our last example is an application to non-survey data. This data set is a geographic one that focuses on the 50 United States and records 12 characteristics for each one, including their position on a range of hot-button issues such as abortion access, gun laws, and the death penalty, as well as more neutral variables such as weather, educational attainment, and whether the state has a professional American football team. The full list of variables is given in Table 2.
On this comparatively small data set our method runs quickly, completing 25 000 Monte Carlo sweeps in under a second and finding three clear classes, as shown in Fig. 7a. The classes correspond to recognized political divisions among the states: there is a group of left-leaning states (so-called “blue states”), a group of right-leaning ones (“red states”), and a group whose positions on the issues are a mix of left and right. Figure 7b shows the division on a map.
Table 2: Summary of the 12 variables used in the 50-states data set. Further details are given in Appendix D, Section D.5.
@@ 4.3.4 Geographic and socio-political traits of the 50 United States
College education Fraction of population with a bachelor’s degree
@@
Gun laws Firearm carry laws
Medicaid Medicaid expansion under the ACA
@@
Same-sex marriage Whether state law permits same-sex marriage
@@
Vote 2024 Winner of the presidential vote in 2024
@@
This data set provides an opportunity to showcase another application of our method, to variable selection—the identification of variables that are, or are not, particularly informative about class membership [49, 11, 18, 62]. In the present case, for instance, where the classes appear to correspond to political persuasion, one might imagine that the “Temperature” variable, which measures annual average temperature, would not be particularly informative. There are a number of methods for performing variable selection in mixture models. One approach that tackles the problem head-on is to calculate the mutual information between the data and the component assignments found by the algorithm [52]. Mutual information is precisely a measure of how much one random variable tells us about another, so it directly addresses the question of how much the responses tell us about class membership. For two general random variables $x$ and $y$, the mutual information $I$ is defined by
@@
where $P(x)$, $P(y)$, and $P(x,y)$ are the marginal and joint distributions of the variables. In the present context, where the variables are the responses $x$ to a question $q$ and assignments to classes $r$, we can calculate these distributions from
@@
5 Conclusions
In this paper we have studied the use of Bayesian mixture models in model-based data clustering and described a new Monte Carlo algorithm for sampling from the integrated posterior of such models, which directly samples component assignments and the number of components, providing a way to perform both clustering and model selection in a single calculation, without the need for additional model selection steps. Our algorithm, a form of collapsed Gibbs sampler that employs rejection-free sampling from the prior over component assignments, is substantially faster than competing algorithms in head-to-head tests, by a factor of up to fifty or more in some cases.
We have demonstrated our approach with applications to Gaussian, Poisson, and categorical models on various data sets, including extensive tests on synthetic data and a selection of real-world examples. We find that the method is consistently able to infer the correct number of classes in the synthetic tests and provides useful results for a range of other tasks, including class assignment, consensus clustering, and variable selection.
[…]
Acknowledgments The author thanks Elizabeth Bruch, Carrie Ferrario, Max Jerdee, Alec Kirkley, and two anonymous referees for useful comments and feedback, and Max Jerdee, Conrad Kosowski, Gabriela Fernandes Martins, and Chethan Prakash for providing the candy dispenser data used in Section 4.2. This work was supported in part by the US National Science Foundation under grant DMS–2404617. Data and code are available online at https://umich.edu/˜mejn/mixture, except for previously published data, which are available at the links and references given in the text and appendices.
Appendix A Proof of correctness for the
Monte Carlo algorithm
In this appendix we prove that the distribution of samples generated by the algorithm of Section 3 converges to the integrated posterior distribution of the corresponding mixture model.